OpenStudy (anonymous):

Use the substitution method to solve the system of equations. x + 2y = 15 2x – 3y = 2 Use the substitution method to solve the system of equations. x + 2y = 15 2x – 3y = 2 @Mathematics

OpenStudy (anonymous):

ok so using that method you take one of your equations and solve for the variable. x=15-2y then you plug that into you next equation and solve for your y 2(15-2y)-3y=2 30-4y-3y=2 -7y=-28 y=4 plug 4 back into the second equation and solve for x then you ahve your answer

OpenStudy (anonymous):

that help?

OpenStudy (anonymous):

mm kinda

OpenStudy (anonymous):

ok what do you need help on?

OpenStudy (anonymous):

idk where you got x from

OpenStudy (anonymous):

the x=15-4y?

OpenStudy (anonymous):

2,4?

OpenStudy (anonymous):